OpenStudy (anonymous):

Without using your calculator, decide which of these is larger: log8 62 or log7 50. How confident are you of your answer? Explain your thinking

OpenStudy (ikram002p):

log8 62 = x1 8^x1= 62 8^2=64 >62 so x1<2.....1 log7 50=x2 7^x2=50 7^2=49<50 so x2>2......1 nw u got x1<2 2<x2 x1<2<x2 x1<x2 log8 62<log7 50

OpenStudy (wolf1728):

In base 8, the log of 8 = 1 and log 8 of 64 = 2 so the base 8 log of 62 is slightly less than 2 In base 7, log 7 = 1 and base 7 log of 49 = 2 the base 7 log of 50 is slightly greater than 2. So, the base 7 log of 49 is slightly larger than the base 8 log of 62. and if you really need to calculate this- here is a good calculator: http://1728.org/logrithm.htm